OpenStudy (anonymous):

(1/(r+2))+(1/r)=((6r-6)/(r^2+2r)) Solve completely check for extraneous solutions

OpenStudy (anonymous):

\[\frac{ 1 }{ r+2 }+\frac{ 1 }{ r }=\frac{ 6r-6 }{ r^2+2r }\]

OpenStudy (anonymous):

ok start by finding the common denominator for the left hand side of the equation

OpenStudy (anonymous):

it be r+2 but how do you get there?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

its LCM is product of r+2 and r

OpenStudy (anonymous):

wait woudn't it be r^2+2r?????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would you have to multiply the top of the fraction too?

OpenStudy (anonymous):

yes you multiply the top and the bottom by what you used to get the common denominator

OpenStudy (anonymous):

\[\frac{ 2r+2 }{ r^2+2r } \right?\]

OpenStudy (anonymous):

yes now look at the denominators on both sides

OpenStudy (anonymous):

thier the same

OpenStudy (anonymous):

Yep so how do you get rid of them from their

OpenStudy (anonymous):

multiply each side by it right?

OpenStudy (anonymous):

yep and they are gone

OpenStudy (anonymous):

and then just solve! god if only my math teacher explained it like this!!!

OpenStudy (anonymous):

don't forget to check for the extraneous solutions

OpenStudy (anonymous):

in this case you can just plug your answer back into the equation and see that it works or look at the original equation and see that \[\neq-2 or 0\] because you can not have 0 in the bottom of the fractions